OpenStudy (anonymous):

Easy yet Hard Question 2+2= (it's not what you think)

OpenStudy (anonymous):

Sure it is: 4 according to mathematical theory. :P

OpenStudy (anonymous):

either i say 4 or 22 you'll sy it's wrong

OpenStudy (saifoo.khan):

\[\Huge 2+2 = 22\]

OpenStudy (anonymous):

@Sjano not a mathematical theory but a fact

OpenStudy (anonymous):

@saifoo.khan: You should seriously go back to grade school :P

OpenStudy (anonymous):

@Zeerak Facts are just theories. Like Math, Evolution and so on. Facts gets altered as the theories get altered .

OpenStudy (anonymous):

You guys are completly missing it the answer is in the question...

OpenStudy (anonymous):

2+2=fish

OpenStudy (anonymous):

\[\left| 22-2 \right|\neq 2 \] \[\forall \epsilon >0\]

OpenStudy (anonymous):

2+2= (it's not what you think) THINK ABOUT IT

OpenStudy (anonymous):

2+2=1+3

OpenStudy (anonymous):

tell the answer already...

OpenStudy (anonymous):

It's not what you think that's the answer

OpenStudy (anonymous):

it equals omega